OpenStudy (asad786):

WILL MEDALWhich expression is equivalent to this one? k · 6 + k · s A. 6(k + s) B. s(6 + k) C. k(6 + s) D. k · (6 + k) · s

OpenStudy (igreen):

Okay, 'k' is being multiplied to '6' and 's'..so what do you think?

OpenStudy (asad786):

I THINK IT A

OpenStudy (igreen):

No..that would be \(\sf 6(k + s) \rightarrow 6 \cdot k + 6 \cdot s\)

OpenStudy (asad786):

isn't that the same thing

OpenStudy (igreen):

What should go on the outside of the parenthesis? Which one is showed twice in \(k \cdot 6 + k \cdot s\)

OpenStudy (igreen):

No..it isn't

OpenStudy (asad786):

oh srry about that

OpenStudy (igreen):

Which is showed twice in \(\sf k \cdot 6 + k \cdot s\)?

OpenStudy (asad786):

c

OpenStudy (igreen):

Yep, you got it.

OpenStudy (asad786):

thanks again i have a few more

OpenStudy (igreen):

\(\sf k(6 + s) \rightarrow k \cdot 6 + k \cdot s\)

OpenStudy (igreen):

No problem, just keep asking.

OpenStudy (asad786):

ut ill open another thread